OpenStudy (anonymous):

Choose the equation below that represents the line that passes through the point (-2, 6) and has a slope of 4. A. y - 2 = 4(x + 6) B. y + 6 = 4(x - 2) C. y + 2 = 4(x - 6) D. y - 6 = 4(x + 2)

OpenStudy (anonymous):

HINT: Rewrite each one in the form: y = mx + b

OpenStudy (anonymous):

Then, you can use process of elimination like this: find which one has slope 4. If there are two or more that have m = 4, then try plugging in (-2,6) into the equations and seeing if LHS = RHS

OpenStudy (anonymous):

There is no need of rewriting all the equations..

OpenStudy (anonymous):

you can just substitute (-2,6) to the equations. :)

OpenStudy (anonymous):

the ansewr is d rite

OpenStudy (anonymous):

You have m = 4 \((x_1, y_1) = (-2, 6)\) use the following equation here: \[\large \color{green}{y - y_1 = m(x - x_1)}\] Put all the values and find the equation..

OpenStudy (anonymous):

@marktierra that is great.

OpenStudy (anonymous):

Just plug in the point and see if LHS = RHS. By process of elimination, you will find one that is correct

OpenStudy (anonymous):

Yes you are right..

OpenStudy (anonymous):

thank u